​(Related to Checkpoint​ 5.4)  ​(Present-value comparison)  You are offered ​$100,000 today or ​$340,000 in 14 years. Assuming that you can earn 16 percent on your​ money, which should you​ choose?If you are offered ​$340,000 in 14 years and you can earn 16 percent on your​ money, what is the present value of ​$340,000​? ​$nothing  ​(Round to the nearest​ cent.)

​(Related to Checkpoint​ 5.4)  ​(Present-value comparison)  You are offered
Answer #1

Present value of ​$340,000=340,000*Present value of discounting factor(rate%,time period)

=340,000/1.16^14

=340,000*0.125195336

=$42,566.41(Approx)

Hence $100,000 today is better having higher present value.
Answer #2

SOLUTION :


It requires to compare the present values of options to select the better one. Higher the PV, better is the option. 


Accordingly, we determine PVs of the two options .


Option 1 :


$100000 today .

Its PV = $100000 .


Option 2 :


Receiving $340000 in 14 years from now. 

Required rate, r = 16% = 0.16

=> 1 + r = 1.16


PV of Option 2 

= FV/(1 + r)^n 

= 340000 / 1.16^14 

= 42566.41 ($) 


PV of Option 2 is = $42566.41 (ANSWER)


So, PV of Option 1 is higher than that of Option 2. 

Hence, Option 1 should be accepted. (ANSWER).



answered by: Tulsiram Garg